- commissioning of 2nd energy BPM (BPM12), also at 500 mm dispersion, but at opposite beta function as BPM24
- commissioning and first data with 4 chicane magnets
- rearranged zygo to BPM4 at center of chicane and BPM3 in front of chicane
- commissioned new UK BPM processor electronics, for latter part of run on BPMs y9->y11 and q10 (split)
- commissioned calibration tone for gain monitoring
- commissioned new calibration method using helmholtz coils with cycles of fast dithering,
preceeded by a BPM4 mover calibration and an energy setpoint calibration afterwards,
automatic processing software also commissioned
- SLAC eLog mirrored to UCL, available under :
- List with runs taken during march '07 run :
- Analysis plans
Priority to reprocess asap the data with 2nd ddc filter implemented and decent calibrations using Helmholtz
macro (esaCalib) : all data is copied to UCL, setting up latest code version.
Note: code changes : ntp3 now calibration tone data ! Higher number ntuples shift.
- Calibration using Dithering and Helmholtz Coils
- Correct step finding algoritm for "pedestal" drift
- restructure esaCalib and put in scale correction with mover calibration and zygo
- re-analyse stability of scales & IQ phases, extend plots like
- Zygo data
- Need to have a look to long term behaviour of the interferometer : plot drifts and vibration modes
- Note: zygo reintialises between runs, so correlate by taking average of beginning of run and offsetting the data
- Magnet data
- sergey has put everything together and obtained the correct beam energy of ~28.5 GeV : continue studies, updated
data files etc..
- energy resolution, compare with BPM12 and BPM24
- comparison with simulation ( sergey & bino )
- compare dispersion in front of & behind chicane... if possible given resolution on 9,10,11
- stability of energy measurement & correlate drifts
- stabiltiy of NMR/Hall probe measurements, study magnet current scans
- Calibration tone data
- Yury implemented an event type flag in the data (ntp2.evtType), which corresponds to real event data, pedestal data or
calib tone data and ntp3 which holds the processed calib info
- study ripple in DDC level, investigate with simple simulation, see https://www.hep.ucl.ac.uk/elog//Bino's+logBook/19
- stability & drifts over time, correlate with temp
- UK BPM Electronics
- check monopole leakage, compare with simulation
- study the FFT of the DDC, gauge the 2omega component leakage when using the old slac cavities
and the 22 MHz IF
- linearity : study q10UKAmp versus q10Amp etc..
- New cavity + mover system
- Status of position & reference cavity : new pictures !!
- Status of mover system
- T474 paper : discuss status
- Contribution to PAC'07 : poster, need author list + update website
- LCWS meeting in Hamburg ( 30 May - 04 June 2007) : contribution ?
- Energy spectrometer meeting in DESY/Zeuthen (06 - 09 June 2007) after LCWS in Hamburg